OpenStudy (anonymous):

Find the marginal cost, marginal revenue, and marginal profit functions. C(x) = 2x; R(x) = 5x − 0.001x2

OpenStudy (rulnick):

marginal cost: C'(x) = 2 marginal revenue: R'(x) = 5 - 0.002x marginal profit (call it P'): P'(x) = R'(x) - C'(x) = 3 - 0.002x

OpenStudy (anonymous):

Find all values of x for which the marginal profit is zero. (Enter your answers as a comma-separated list.) x =

OpenStudy (rulnick):

Just set the marginal profit value to zero and solve for x: 0 = 3 - 0.002x 0.002x = 3 x = 1500

OpenStudy (anonymous):

Thank you.